OpenStudy (anonymous):

z -1 5 ----- = ----- z+4 7

OpenStudy (anonymous):

solve for z

OpenStudy (anonymous):

cross multiplying gives (z-1)7 = 5(z + 4) 7z -7 = 5z +20 7z - 2z =20 + 7 2z = 27 z = 27/2

OpenStudy (anonymous):

thanks again